OpenStudy (anonymous):

Find an approximate solution to the equation e^(x-9)=1/10

OpenStudy (mertsj):

Did you take the natural log of both sides?

OpenStudy (nurali):

e^(x-9)=1/10 ln both sides ln e^(x-9)=ln1/10 (x-9)lne=-2.3 lne=1 x-9=-2.3 x=-2.3+9 x=6.7